June 2, 2016 Calorimeter

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Preparation for REST production
  4. Calibrations
  5. Reconstruction and Simulation
  6. Any other business


Attendees:Elton, Jon, Adesh (JLab); Curtis, Will (CMU); Christina, George (Athens); Sean (NW); Matt (IU); Zisis, Ahmed, Tegan, Colleen, Andrei (UofR)

  1. Announcements
  2. Action Items
  3. Preparation for REST production
    1. FCAL (Ahmed)
      • Pedestals, gain constants, non-linear corrections to be checked in today
      • Non-linear corrections use a piece-wise linear and exponential forms. Will show this next time.
      • CCDB does not presently have sufficient entries for all constants, some are hard coded for now. Later updates will place everything into the db.
      • Also need to check in code to match the new correction function.
      • Expect to have everything in place by this evening. Possible tweaking of constants may happen tonight. Adesh will keep Sean informed of progress.
      • Pluggin to plot pi0 peaks with various cuts has been checked in.
      • (Sean) Launch will run off a fixed time stamp in the ccdb.
  4. Calibrations
    1. BCAL (Will)
      • Gain constants are checked in for this launch
      • Non-linear corrections are being studied by looking at changes of the pi0 mass vs various quantities, but are not included at this time. Changes of pi0 mass are less than 2 MeV over range of parameters studied.
    2. BCAL Timing
      • George: No new update since updating code to match new calibrations for zpoint. After discussions with Zisis and Christina, he decided to move to work on physics analysis and avoid overlaps with Andrei's work.
      • Andrei: Time walk and time offset calibrations have been completed for charged particles.
      • There is no time to get them ready for the launch later today
      • Calibration with charged tracks shows clear tof bands for protons and pions. Checking systematics as a function of layer number and other parameters.
      • Starting analysis with neutral showers.
    3. BCAL LED Monitoring (Ahmed)
      • https://logbooks.jlab.org/entry/3406797
      • Good progress in determining different Led configurations, but some anomalies still remain, e.g. in column 1 in logbook plots
      • Sean has written the skim program to peel off the LED triggers. In addition to stripping out events with the front panel (FP) trigger bit 9 or 10 set, events are selected for a)Total energy in BCAL > 12 GeV or b) Number of hits in BCAL > 200.
  5. Reconstruction and Simulation
      • Colleen and Tegan are looking at Michel electrons from muon decay.
      • Colleen wrote a toy MC to look at energy depositions.
      • This toy MC is now being used as an event generator to seed hdgeant with the energy and verteces of electrons.
      • Comparisons between the toy MC and hdgeant are underway.
    1. Fcal efficiency (Jon)
      • Has been investigating the use of omega->pi0 gamma to determine Fcal efficiencies.
      • The fully exclusive channel including all photons works reasonably well.
      • Excluding the bachelor photon from the reconstruction leads to a S/B ~ 1/50, which will not allow for the efficiency to be determined with much accuracy.
      • Started looking at omega -> pi+pi-pi0 and this channel is more promising.
  6. Any other business
    1. Next launch (Adesh)
      • Second launch is about 2 weeks away. The data set comes from the last weekend of the run period.
      • Need to start getting calibrations ready for the new launch
      • Adesh discovered that DAQ values were changed affecting pedestals for about 10-15 runs. Special calibrations of pedestals, gains and possibly non-linear corrections will be needed for these runs. (Changes were not noted in the log book).
      • We need to have better control over DAQ adjustments during the run period.